satellite73:

\[\frac{dy}{dx}-y=6\] is linear, you can look up what that means

Vrefela:

ok

satellite73:

idea is this

satellite73:

you want to make the left hand side look like the derivative of a product, my multiplying by the "integrating factor"

Vrefela:

Alright

satellite73:

in other words you find something, the book calls it \(\mu\) so that when you multiply you will have \[\mu \frac{dy}{dx}-\mu y=6\mu\] adn \(\mu=\mu(x)\) a function of x

satellite73:

damn typo

satellite73:

i'll get this eventually

satellite73:

example in the book is \[\frac{dy}{dx}-3y=6\]

satellite73:

so now how to find \(\mu\) looks complicated but it is easy

Vrefela:

Looks tough honestly

satellite73:

first we have \(-3\) in front of the y step one is to integrate wrt x i.e .\[\int-3dx\]

satellite73:

let me know what you get

Vrefela:

Ok -3x

satellite73:

right step two is take \(e^{-3x}\) and we rare done, that is \(\mu\)

Vrefela:

Why?

Vrefela:

And we are looking for mu?

satellite73:

lets see why

satellite73:

we found \(\mu\) it is \(e^{-3x}\)

Vrefela:

Thats it?

satellite73:

ok now lets multiply and see what magic happens

satellite73:

oh no, that i just \(\mu\) we are not done lets multiply both sides by \(e^{-3x}\)

satellite73:

\[\frac{dy}{dx}-3y=6\] \[e^{-3x}\frac{dy}{dx}-3e^{-3x}y=6e^{-3x}\]

satellite73:

now look at the left hand side and see what you have why, it is the derivative of \(e^{-3x}y\) a miracle check it

Vrefela:

one sec brb

satellite73:

by the produce rule \[\frac{d}{dx}[e^{-3x}y]=e^{-3x}\frac{dy}{dx}-3e^{-3x}y\]

Vrefela:

back

Vrefela:

Ok ?

satellite73:

ok

satellite73:

so is it clear that the left hand side is a derivative, the derivative of \[e^{-3x}y\]?

Vrefela:

yes

satellite73:

then to solve this \[e^{-3x}\frac{dy}{dx}-3e^{-3x}y=6e^{-3x}\] integrating you get \[e^{-3x}y=\int 6e^{-3x}dx+c\]

satellite73:

let me know when that is clear

Vrefela:

So -2e^(-3x) + c?

satellite73:

final answer \[e^{-3x}y=-2e^{-3x}+c\]

Vrefela:

Okie

satellite73:

actually final answer is best to solve for y

Vrefela:

So divide by e^(-3x)

satellite73:

multiply both sides by \(e^{-3x}\) get \[y=-2+ce^{3x}\]

Vrefela:

Oh right right

satellite73:

oops i meant multiply both sides by \(e^{3x}\)

satellite73:

check that is works

satellite73:

want to try one more?

Vrefela:

Please

satellite73:

ok let me pick one first step is to find \(\mu\) we can do say 5

Vrefela:

Okie dokie